Dear team,
Please could you Help me with the task below!
The sum of abcde and edcba is a five-digit number and is a multiple of 11111
Also the number abcde is a multiple of seven and the sum of its digits is 10
(a,b,c,d and e are all different digits)
Find the number abcde
And then using ANY mathematical operations (such as +; -; *; /; Log; etc.)
between the digits abcde in THIS sequence get the result to be 20
I need to send this by Sunday 19 december if possible!
Hoping this is not to much trouble.
Thank you.
Kindest wishes
Nasy

Answer by Edwin McCravy(20054)

abcde = 14203
edcba = 30241
-------------
  SUM = 44444 = 4×11111

a+b+c+d+e = 1+4+2+0+3 = 10

1×4×(2+0+3) = 4×5 = 20

Edwin